# MMICT: Boosting Multi-Modal Fine-Tuning with In-Context Examples

### 11 Dec 2023 v1

@@ 1 Introduction
There are three mainstream methods that integrate the visual encoder and LLM in MM-LLM: 1) The addition of layers and utilization of cross-attention in the LLM, exemplified by Flamingo and Otter, which incurs high computational costs. 2) The incorporation of learnable layers subsequent to the visual encoder, facilitating the conversion of features for language model comprehension, as demonstrated in the BLIP2 (Li et al. 2023c) and MiniGPT4 (Zhu et al. 2023), which offers a simpler and less hardware-intensive approach. 3) The inclusion of supplementary learnable parameters in the LLM, as seen in the LLaMA-Adapter (Gao et al. 2023), which yields a faster training speed albeit with marginally lower performance in comparison to alternative approaches. Considering the pros and cons of the aforementioned methods, we opt for BLIP2 as our baseline and propose MMICT.
While building upon BLIP2, MMICT represents a significant advancement by introducing a novel dimension. BLIP2 adopts traditional fine-tuning strategy with only query inputs, and exclusively utilizes the cross-modal pre-trained Qformer to extract visual features. To capture the above-mentioned multi-modal features within a unified model architecture, we design the Multi-Modal Hub (M-Hub) used in MMICT. Different from Qformer, M-Hub can produce either uni-modal features or multi-modal features that fuse information from different modalities. Considering that different demonstration factors (e.g., feature extraction strategy, sampling number for demonstrations and sampling strategies for demonstrations) may heavily affect the performance, we design various variants of in-context demonstrations by leveraging the flexibility of M-Hub. In summary, the contributions of this work are:
@@ 3.4 Inference
Table 1: Instruction templates used for image captioning and video captioning.
@@ 4.9 The Impacts of Feature Levels
4.10 Direct Comparison with BLIP2
@@
To further demonstrate the advantages of MMICT, we provide a direct comparison with BLIP2 in Tab. 8. As displayed in Tab. 8, the performance of MMICT shows a remarkable improvement when the image encoder is further fine-tuned, as opposed to just fine-tuning the M-Hub on the COCO dataset. Additionally, MMICT still outperforms BLIP2 that we re-implemented under their configuration. Significantly, the performance gap between MMICT and BLIP2 becomes even larger, indicating that the benefits of MMICT amplify as more parameters are trained.

### 12 Dec 2023 v2

@@ 1 Introduction
MMICT is built based on BLIP-2 (Li et al. 2023c). BLIP-2 adopts a traditional fine-tuning strategy with only query inputs, and exclusively utilizes the cross-modal pre-trained Qformer to extract visual features. To better capture the multi-modal features within a unified model architecture, we design the Multi-Modal Hub (M-Hub) used in MMICT. Different from Qformer, M-Hub can produce either uni-modal features or multi-modal features that fuse information from different modalities. Considering that different demonstration factors (e.g., feature extraction strategy, sampling number for demonstrations and sampling strategies for demonstrations) may heavily affect the performance, we design various variants of in-context demonstrations by leveraging the flexibility of M-Hub.
In summary, the contributions of this work are:
@@ 2.1 Multi-Modal Large Language Models
In summary, there are three mainstream methods that combine visual encoder and LLM into MM-LLM: 1) The addition of layers and utilization of cross-attention in the LLM, exemplified by Flamingo and Otter, which incurs high computational costs. 2) The incorporation of learnable layers subsequent to the visual encoder, facilitating the conversion of features to enhance the comprehension of LLM, as demonstrated in the BLIP-2 and MiniGPT4 (Zhu et al. 2023), which offers a simpler and less hardware-intensive way. 3) The inclusion of supplementary learnable parameters in the LLM, as seen in the LLaMA-Adapter (Gao et al. 2023), yields a faster training speed albeit with marginally lower performance in comparison to alternative approaches. Considering the pros and cons of the aforementioned methods, we opt for BLIP-2 as the base of MMICT.
@@ 3.4 Inference
Table 1: Instruction templates used for captioning tasks.
@@ 4.9 The Impacts of Feature Levels
4.10 Direct Comparison with BLIP-2
@@
To further demonstrate the advantages of MMICT, we provide a direct comparison with BLIP-2 in Tab. 8. As displayed in Tab. 8, the performance of MMICT shows a remarkable improvement when the image encoder is further fine-tuned, as opposed to just fine-tuning the M-Hub on the COCO dataset. Additionally, MMICT still outperforms BLIP-2 that we re-implemented under their configuration. Significantly, the performance gap between MMICT and BLIP-2 becomes even larger, indicating that the benefits of MMICT amplify as more parameters are trained.